OpenStudy (anonymous):

Find 3 Ordered Pairs: 3x+2y=2

OpenStudy (anonymous):

choose number like 0 1 and 2 for x

OpenStudy (anonymous):

3(0)+2(1)=2 Would This Work

OpenStudy (anonymous):

3x - 2y)(3x - 2y) = 9x² - 6xy - 6xy + 4y² = Collect like terms 9x² - 12xy + 4y²